MONT 104Q – Mathematical Journeys: From Known to Unknown Some Questions to Think About Concerning the Odyssey

1. The narrative of the Odyssey is not a simple recounting of Odysseus’s journey home in chronological order. For one thing, the poem starts right in the middle of the story. How does this structure work? Who is “on stage” for most of the first few books? Where and how do we meet the main character, Odysseus? How is his story different from those of the other Greek heroes who returned from Troy (Agamemnon, Menelaus, Nestor, and the rest)? Where and how does most of Homer’s story take place, and how is the background filled in? What effects does this have on the hearer/reader?

2. We often seem to think that, to be exciting, a story requires surprise and suspense. Yet the most familiar episodes of the Odyssey where Odysseus has been in real peril appear as “flashbacks” during Odysseus’ recounting of his travels in the Phaeacian court of Alcinous in Books 9 - 12 (and so we know beforehand that he has survived them). Even the eventual conclusion of the Odyssey is foretold multiple times – as a prophecy by the seer Tiresias, by the ghost of Agamemnon during Odysseus’ visit to Hades recounted in Book 11 (at least conditionally), and even by Penelope herself in her recounting of her dream of the geese killed by the eagle (in Book 19). How and why are Homer’s storytelling techniques different from those with which we are more familiar?

3. What are Odysseus’s key character traits? How are these brought out in episodes like the encounter with the cyclops Polyphemus? For the Greeks portrayed in the Odyssey, do you suppose being called “tricky, crafty” or “wily” was a compliment or an insult? (And look at Athena’s reaction to Odysseus’ initial lies when he meets her in Book 13.)

4. What do we know about the meaning of Odysseus’ name, as explained by his grandfather Autolycus? (See Book 19, lines 440 - 452 of Lombardo’s version.)

5. How successful would you say Odysseus is as a military leader, based on what happens and what we learn during the Odyssey? Is it his fault that all of his men are killed on the return trip? How well does he lead his most trusted men like Eurylochus in the episodes in Book 12 (the Sirens, Scylla and Charybdis, the cattle of Helios)? And does a good leader continually blame his men for his problems as Odysseus often does?

6. Odysseus is not the only character who undergoes a journey in this story. In what sense can we say that his son Telemachus also has this sort of experience? Where does he start; what famous characters from the story of the Trojan war does he meet; where does he end? What changes does he undergo?

7. For modern readers, one possibly strange aspect of the world of the Odyssey is the extent to which the Olympian gods (Zeus, Hera, Athena, Apollo, Artemis, Aphrodite, Ares, Poseidon, Hermes, etc.) are seen as taking a direct hand in the events of the story through interventions in favor of, or against, mortal characters. In this world, is there any room for humans to exercise free will? Can the choices humans make make a difference? What kind of religion do the characters in the Odyssey seem to follow? How are the Greek gods described – not just their appearance, but also their “personalities?” What sorts of relationships do the characters in the Odyssey have with the gods?

8. What does justice mean for the people in the Odyssey? Are revenge and retribution the only options, or are there other kinds of justice? (For instance, think about the ways Odysseus treats different servants in his house on his return.)

9. If you were away fighting a war for 10 years and then took another 10 years to travel home, hopping into bed with the minor goddesses Circe and Calypso along the way (and perhaps also encouraging the Phaeacian princess Nausicaa’s unrequited crush on you), would it be reasonable for you to expect a wife to wait faithfully for your return? Wouldn’t it be natural for her to want to get on with her own life?

10. Penelope has often been seen as the perfect embodiment of chastity and marital fidelity, but is there another motivation for her actions that is briefly hinted at? What does this say about Penelope’s character, and/or the kinds of roles women (and goddesses) play in the society depicted in the Odyssey? What sorts of power can women (or goddesses) hold and why?

11. There is a small hint in Book 19 that Penelope recognizes that the old beggar that has arrived is actually her husband Odysseus right from the start. What is that hint, and why do you suppose she doesn’t open up to Odysseus sooner?

12. Isn’t question 9 somewhat unfair to Odysseus? Was he an entirely willing partner for Circe and Calypso, or was there some coercion involved? What about Odysseus’s and Nausicaa’s feelings for one another? (As far as we learn from Homer, their relationship was chaste, but note that Odysseus never mentions her to Penelope when he talks about his experiences.)

13. How would you describe Penelope’s suitors, especially Antinous? Would you want to have anything to do with them? But, on the other hand, think about this: By the chronology of the story, they must actually be the children of the men who accompanied Odysseus to Troy and who have all been killed on the return trip(!) Does this change how you think about them? Is it possible for a modern reader to accept the way Odysseus deals with them? If the events described here happened in real life today, wouldn’t Odysseus be treated as a mass murderer? How have our values changed from those depicted in the Odyssey? Does this change in values affect your enjoyment of Homer’s poem?

14. People in the Odyssey have no problem boasting, but at times they also warn against excessive pride. Where is the dividing line? When is boasting OK, and when is humility required?

15. What do think accounts for the fact that the Odyssey has been such a big part of our literature for such a long time (i.e. consistently since it was first written down about 800 BCE)? What do people find so appealing about it and why has it inspired modern satirical reworkings like James Joyce’s novel Ulysses or the Coen brothers’ film Oh Brother, Where Art Thou?

16. Before they were written down, the Greek epic poems such as the Odyssey (and the Iliad, its surviving companion piece also ascribed to Homer) were almost certainly performed orally with music by bards who knew the whole poems and recited sections from memory for audiences. To get a feeling for what this might have been like, look again at the description of the songs of Demodocus in Book 8 (and Odysseus’ high praise for them: “When you sing about the fate of the Greeks who fought at Troy, you have it right, all that they did and suffered, all they endured. It’s as if you had been there yourself, or heard a first-hand account.”) What traces of that history can we see in the surviving text, and how has our translator, Stanley Lombardo, chosen to render some of those features in the version we are reading?